What is the future value (in $) of cash flows 1-3 at the end of year 3, assuming a 6% interest rate (compounded annually)? End of year Cash flow 1 $600 2 829 3 710 4 3,500 5 1,250 6 4,530 7 2,350
Answer #1

Future Value of cash flow after "n' years is given by = CF*(1+Rf)^n

Where Rf is risk free rate

Now 1st year Cash flow = $600

Hence its Future Value after end of year-3 = 600*(1+.06)^2 = 674.16

Similarly Future Value of Year-2 Cash flow at the end of Year-3 = 829 *(1+.06)^1 = 878.74

And

Cash Flow of Year-3 is received at the end of year hence its future value = 710

So, Future Value of Total Cash flows of Year (1to3) = $2262.90